OpenStudy (anonymous):

If the point (4,-1/3) lies on the graph of the equation 2x+cy=11, find the value of c.

jimthompson5910 (jim_thompson5910):

2x+cy=11 2*4+c(-1/3)=11 ... plug in x = 4 and y = -1/3 now solve for c

OpenStudy (anonymous):

Thanks! I was unsure of how to start this.

OpenStudy (anonymous):

jims right :)

OpenStudy (anonymous):

I got -9.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

np